MOLECULAR BEAM INFRARED SPECTRUM OF THE HF STRETCHING FUNDAMENTAL BAND OF THE HF-BF_{3} COMPLEX

K. NAUTA and R.E. MILLER, Dept. of Chemistry, University of North Carolina, Chapel Hill, NC 27599; G.T. FRASER and \underline{W.J. LAFFERTY}, Optical Technology Division, NIST, Gaithersburg, MD 20899.

The microwave spectrum of the antihydrogen bonded HF-BF_{3} complex has recently been reported by Philips et al. ^{a}. The spectrum reveals a near symmetric F-BF_{3} heavy atom frame with the proton slightly off axis, and a small barrier to internal rotation of the H atom against the F atoms. Here, we have investigated the H-F stretching fundamental band of this complex using a molecular beam optothermal spectrometer and a color-center laser. Two subbands are expected: a parallel band with Δ m = 0 selection rules and a perpendicular band with Δ m = ±1 selection rules. At present, only the perpendicular band has been observed. This band has the appearance of a perpendicular band of a symmetric rotor, but with anomalously strong ^{r}R transitions and weak ^{p}P transitions. Upper state constants are obtained fixing the ground-state constants to the microwave-determined values. The band has a large positive ζ Coriolis constant, from which an estimate of the barrier to internal rotation can be made. A search is presently underway to locate the expected weaker Δ m=0 parallel component.

aJ. A. Philips, M. Canagaratna, H. Goodfriend, A. Grushow, J. Almlof and K. R. Leopold J. Am. Chem. Soc. \underline{117}, 12459 (1995)